OpenStudy (anonymous):

Simplify: (-5+3)^3 - (-2-1)^2

OpenStudy (unklerhaukus):

\[(-5+3)^3 - (-2-1)^2\]\[=(-2)^3-(-3)^2\]

Parth (parthkohli):

\(\Large \color{purple}{\rightarrow (-2)^3 + 3^2 }\) \(\Large \color{purple}{\rightarrow -8 + 9 }\) \(\Large \color{purple}{\rightarrow 1 }\)

Parth (parthkohli):

Do you get it?

OpenStudy (unklerhaukus):

\[=-8−9\]

OpenStudy (anonymous):

answer choices: a. -17 b. 17 c. -15 d. 15

OpenStudy (unklerhaukus):

ParthKohli has made a simple error

OpenStudy (unklerhaukus):

\[(-3)^2=-3\times-3=9\]

OpenStudy (unklerhaukus):

\[-8-9=-17\]